# Evaluating Open Language Models Across Task Types, Application Domains, and Reasoning Types: An In-Depth Experimental Analysis

### v1

@@ Abstract
The rapid rise of Language Models (LMs) has expanded their use in several applications. Yet, due to constraints of model size, associated cost, or proprietary restrictions, utilizing state-of-the-art (SOTA) LLMs is not always feasible. With open, smaller LMs emerging, more applications can leverage their capabilities, but selecting the right LM can be challenging. This work conducts an in-depth experimental analysis of the semantic correctness of outputs of 10 smaller, open LMs across three aspects: task types, application domains and reasoning types, using diverse prompt styles. We demonstrate that most effective models and prompt styles vary depending on the specific requirements. Our analysis provides a comparative assessment of LMs and prompt styles using a proposed three-tier schema of aspects for their strategic selection based on use-case and other constraints. We also show that if utilized appropriately, these LMs can compete with, and sometimes outperform, SOTA LLMs like DeepSeek-v2, GPT-3.5-Turbo, and GPT-4o.  [^1] GitHub repository containing the code implementation of this work: https://github.com/neelabhsinha/lm-application-eval-kit
Evaluating Open Language Models Across Task Types, Application Domains, and Reasoning Types: An In-Depth Experimental Analysis
@@ 1 Introduction
The field of NLP has advanced significantly with the development of Language Models (LMs). Accelerated development of LMs Brown et al. (2020); Touvron et al. (2023); Almazrouei et al. (2023); Team et al. (2024); Abdin et al. (2024); DeepSeek-AI (2024) has expanded their application across numerous types like Title Generation (Keles and Bayraklı, 2024), Data Exploration (Ma et al., 2023), Dialogue act recognition (Qiang et al., 2024); domains like Economics & Finance (Rajpoot et al., 2024; Yu et al., 2023), Politics (Feng et al., 2023; Kuila and Sarkar, 2024), Nutrition & Food (Yang et al., 2024), News (Kuila and Sarkar, 2024); and reasoning types (Huang and Chang, 2023) like ANALOGICAL (Wijesiriwardene et al., 2023) and Multi-hop (Pan et al., 2021) reasoning.
@@
Initially, LMs were relatively weak like GPT-2 (Radford et al., 2019), too large in size like GPT-3 (Brown et al., 2020), expensive like GPT-4 (OpenAI et al., 2024), and/or closed and accessible only via APIs. This posed different challenges of accessibility, computational resource limitations (Ding et al., 2024), costs (Jimenez Gutierrez et al., 2022) (GPT-4 Turbo charges \$10/1M input, \$30/1M output tokens), data privacy, inflexibility of customization, and so on. However, there has been recent rise in competitive LMs which are relatively small and openly available. These properties remove the aforementioned limitations, and provide additional benefits like on-device usage, faster inference time, easier compliance and security management.
@@
But while large SOTA LMs universally perform well out-of-the-box, smaller LMs may not fit every use case due to variation in training data, pre-training strategies, and architectural decisions. They can have their own strengths, weaknesses and may work differently with different prompts. Further, utilization strategies like zero-shot usage, customizing pre-trained models (e.g., fine-tuning (Mosbach et al., 2023)), using in-context learning (Wei et al., 2022a; Dong et al., 2023), prompt engineering (Brown et al., 2020) can differ with varying need. Writing helpful prompts also requires time and domain expertise. Conducting all this analysis with constraints of time, money, computational resources is a complicated task. Technical reports of some LMs (Team et al., 2024; Abdin et al., 2024) report performance on different benchmarks, but they are not elaborate. So, there is a need for a comprehensive evaluation of LMs across different task types, application domains, reasoning types, etc., which can provide insights of their capabilities for different use-cases.
@@
To bridge this gap, we conduct an in-depth experimental analysis of performance of 10 openly available LMs between 2B–11B parameters using diverse prompt styles like chat-style task definitions, in-context examples on three aspects – task types, application domains and reasoning types, which can be considered as three different ways of categorizing an application area aiming to use them. We first build a schema by selecting 12, 12, and 10 entities from each aspect respectively in English language from Super-Natural Instructions (Wang et al., 2022), and group similar entities together. For example, under application domains, we group ‘Social Media’ and ‘News’ in ‘Media and Entertainment’. This three-tier structure (aspect, group, entity) will allow finding patterns in capabilities of LMs at multiple level, along different aspects.
@@
A recent work (Zhao et al., 2024) dives into efficacy of LoRA (Hu et al., 2022a) fine-tuning of smaller LMs, but keeps prompt as a constant, and limits to varying task types only. Further, usage strategy of LMs can significantly vary – fine-tuning (Mosbach et al., 2023), PEFT (Han et al., 2024), multi-stage pipeline based, or directly with/without prompt engineering due of lack of sufficient data, resources, or need for customization. Hence, we consider semantic correctness of generated outputs as a measure of LMs’ innate ability to perform a task, and evaluate five pre-trained and five instruction-tuned (IT) (Ouyang et al., 2022) LMs out-of-the-box using 8 different prompt styles. The outputs can later be aligned if needed using any desired technique. We also show that some of these LMs can meet most needs, and match or outperform models like GPT-3.5-Turbo (Brown et al., 2020; OpenAI, 2023), GPT-4o (OpenAI et al., 2024; OpenAI, 2023) and DeepSeek-v2 (DeepSeek-AI, 2024). Our analysis aims to identify the best LM and prompt style for a given requirement, and also provides relative performance insights of different family of LMs, scale of LMs, and between pre-trained and IT models.
In this work, we try to address the following research questions: (i) What are some advantages and limitations of using small, open LMs over large, proprietary LMs? (ii) Given external system or functional constraints, how do different LMs perform for one’s usage needs? (iii) How do their performance vary with different prompts? (iv) How to effectively use them for best results? (v) Where do they stand compared to large SOTA models?
We make the following key contributions, which are also illustrated in Figure 1:
@@
(i) Propose a three-tier schema for analysis of performance of LMs for different task types, application domains and reasoning types.
@@
(ii) Conduct an in-depth experimental analysis of semantic correctness of outputs of 10 openly available LMs in 2B–11B size across these aspects.
(iii) Analyze and study the performance of LMs at aspect-level with multiple prompt styles.
(iv) Show that some of these LMs can compete with (and sometimes outperform) large SOTA LMs like GPT-3.5-Turbo, GPT-4o and DeepSeek-v2.
2 Experimental Setup
@@
In this section, we describe our experimental setup in detail discussing dataset, prompt styles, LMs, selection process of aspects, and evaluation metrics.
2.1 Dataset for Experiments
@@
We derive our experimental dataset from Super-Natural Instructions (Wang et al., 2022), containing multiple tasks with expert-curated instructions. Each task has a definition describing a task in chat-style, in-context examples, and multiple instances (refer Figure 2 for an example). We also use the given labels of task type describing nature of a task (eg. question answering, data to text, etc.), domain describing the field of the task (eg. history, news, etc.), and reasoning type, describing the type of reasoning (if any) needed in the task (eg. multihop, analogical, etc.) to for categorization.
@@
2.2 Prompts
We conduct our experiments using multiple prompt styles - including/excluding task definitions, and with 0, 2, 4, 8 examples for each instance. Examples help LMs (even pre-trained) with in-context learning (Wei et al., 2022a; Dong et al., 2023) without altering their parameters. This is followed by an actual task instance. We select examples from positive examples section of the task. So, we get total 8 prompt styles for each task instance. An example of prompt with definition and 2 examples is given in Figure 2. ‘Input’ and ‘Output’ are used because they can be universal for all tasks.
2.3 Language Models Used
The focus for this work is on open LMs from 2–11B parameters for adaptability and computational efficiency. We experiment with five pre-trained models and five instruction-tuned (IT) models. Analysis of pre-trained models, trained for next-word prediction, will give an insight into LMs’ ability and knowledge to perform the tasks. They can either be used directly or customized further by fine-tuning or any other desired method. IT models will suit where out-of-the-box performance on chat-style human instructions may be desired due a simple use-case or due to unavailability of sufficient data/resources to customize the models.
Our experiments utilize newly released, open, SOTA small-scale models – Gemma-2B, Gemma-7B (Team et al., 2024), Llama-3-8B Touvron et al. (2023); AI@Meta (2024), Mistral-7B-v0.3 Jiang et al. (2023), and Falcon-2-11B Almazrouei et al. (2023); TIIUAE (2024). We also take their instruction-tuned (IT) (except Falcon-2-11B - not available), and Phi-3-mini-128k-instruct Abdin et al. (2024). In this paper, suffix "-I" indicates instruction-tuned. This gives us a good range of models across sizes, families and pre-trained/IT. Check Appendix F for more implementation details describing experimental settings.
@@ 2.4 Selection of Aspects
We analyze results of LMs across all three aspects – task types, domains and reasoning types. There are many entities in each aspect in the dataset. Therefore, for brevity, we filter and rearrange these to create a schema given in Figure 3. Our intention is to provide a structure to these kinds of study and cover a wide spectrum of entities. Understanding some of the definitions, specifically for reasoning types may be difficult, but is detailed more in a survey (Guo et al., 2023) and the dataset repository [^2] https://instructions.apps.allenai.org.
For schema, we select diverse entities from each aspect and group them. For example, ‘Fiction’ and ‘Books’ are grouped under ‘Art and Literature’. This allows analysis at three levels of hierarchy - aspect, group and entity level, which is how we address them in rest of this paper. Some tasks can overlap between entities of same aspect (Kuila and Sarkar, 2024) or different aspects (Keles and Bayraklı, 2024), and some may not belong to an aspect. There are more entities in domains and reasoning types not included here for brevity but listed and evaluated in Appendix B.1.
@@ 2.5 Evaluation Metrics
Performance is initially reported using ROUGE 1/2/L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore P/R/F1 (Zhang et al., 2019). However, there are different limitations of evaluating LM outputs using these (Grusky, 2023; Tanya Goyal, 2022; Xiao et al., 2023; Sai et al., 2021). Since LMs can be fine-tuned/adapted for alignment (Zhou et al., 2023), we evaluate their knowledge via semantic correctness of outputs using BERTScore recall with roberta-large (Liu et al., 2019), based on insights from recent works (Sai et al., 2021; Xiao et al., 2023).
Some tasks, like classification, aren’t generation tasks. But we still consider them as one. It gives a uniform evaluation paradigm, and by aligning outputs using fine-tuning/ICL (Zhao et al., 2023), verbalizers (Hu et al., 2022b), post-processing, labels can be obtained from language outputs.
@@ 3 Experiments and Results
We use all the prompt styles with each of the task instance, do a forward pass on the LM, and decode the output using greedy decoding, which is evaluated using mean of BERTScore recall with available references. We also experimented with other sampling techniques (Holtzman et al., 2020), but didn’t find any improvement (refer Appendix D). So, we used greedy, due to its reproducible nature. The following subsections describe our experiments, outlines the results and discusses them.
3.1 Overall Performance of LMs
First, we evaluate all LMs across multiple evaluation metrics globally using the best prompt style in Table 1 (relative performance across prompt styles is reported in Appendix E.1).
Model # Params Def Ex. R-1 R-2 R-L METEOR BERTScore P/R/F1
Gemma-2B 2.51B 4 22.04 7.88 21.23 18.12 78.22/86.41/81.88
Mistral-7B 7.25B 8 1.17 0.54 1.08 1.99 49.25/58.41/53.40
Gemma-7B 8.54B 0 18.17 5.89 17.49 16.14 71.86/81.06/75.94
Llama-3-8B 8.03B 0 16.38 5.35 15.30 14.96 75.52/82.73/78.80
Falcon-2-11B 11.1B 8 16.88 6.46 16.01 16.45 79.65/86.18/82.72
Gemma-2B-I 2.51B 2 27.56 8.08 26.24 20.62 84.56/88.06/86.19
Phi-3-mini-128k-I 3.82B × 0 7.17 3.17 6.26 8.05 55.47/60.28/57.72
Mistral-7B-I 7.25B 8 51.96 14.67 50.12 35.55 91.29/93.76/92.39
Gemma-7B-I 8.54B 0 8.64 3.23 7.96 12.57 78.18/85.14/81.48
Llama-3-8B-I 8.03B × 8 4.68 2.19 4.23 8.31 74.23/84.33/78.89
@@ 3.1 Overall Performance of LMs
We also find that the correlation between BERTScore recalls of LMs, shown in Figure 4, is low. This strengthens our hypothesis that different LMs have different strengths. Since task types, domains and reasoning types are key aspects of variations, we analyze the performance on them as per Figure 3. For these analyses, we use the best prompt style for that entity of that aspect (refer Appendix E.2 to determine that).
@@ 3.2 Comparison Across Task Types
(d) Task Types; Instruction-tuned Models
(e) Domains; Instruction-tuned Models
(f) Reasonings; Instruction-tuned Models
@@
Most of the pre-trained models perform reasonable well on most tasks, except is Mistral-7B, which is not able to compete with other models in most tasks, specifically for classification & recognition, and generation tasks. Gemma-2B performs better than all 7B and 8B models. It is the best across 50% of the task types, with Falcon-2-11B leading in the remaining. Considering the scale of the two models, Gemma-2B is a strong choice with resource constraints across all task types, unless Falcon-2-11B is needed purely on performance. Gemma-7B and Llama-3-8B hover below the top two with changing differences.
In IT models, Mistral-7B-I performs best on all task types, with Gemma-2B-I as second-best (except overlap extraction). The difference is smaller for linguistic relationship and generation tasks, but large for semantic & pragmatic analysis tasks. Like their pre-trained variants, Gemma-7B-I and Llama-3-8B-I seldom compete with Gemma-2B-I in some tasks, but never outperforms it. Phi-3-mini-128k-I behaves very sensitive, and only competes in 2 task types. So, Gemma-2B and Mistral-7B-I can be selected on performance and resources trade-offs.

### v2

# Are Small Language Models Ready to Compete with Large Language Models for Practical Applications?

@@ Abstract
The rapid rise of Language Models (LMs) has expanded their use in several applications. Yet, due to constraints of model size, associated cost, or proprietary restrictions, utilizing state-of-the-art (SOTA) LLMs is not always feasible. With open, smaller LMs emerging, more applications can leverage their capabilities, but selecting the right LM can be challenging as smaller LMs don’t perform well universally. This work tries to bridge this gap by proposing a framework to experimentally evaluate small, open LMs in practical settings through measuring semantic correctness of outputs across three practical aspects: task types, application domains and reasoning types, using diverse prompt styles. It also conducts an in-depth comparison of 10 small, open LMs to identify best LM and prompt style depending on specific application requirement using the proposed framework. We also show that if selected appropriately, they can outperform SOTA LLMs like DeepSeek-v2, GPT-4o-mini, Gemini-1.5-Pro, and even compete with GPT-4o.  [^1] GitHub repository containing the code implementation of this work: https://github.com/neelabhsinha/lm-application-eval-kit
Are Small Language Models Ready to Compete with Large Language Models for Practical Applications?
@@ 1 Introduction
The field of NLP has advanced significantly with the rise of Language Models (LMs). Rapid development of LMs Brown et al. (2020); Touvron et al. (2023); Almazrouei et al. (2023); Team et al. (2024b); DeepSeek-AI (2024) has expanded their use across numerous types like Title Generation (Keles and Bayraklı, 2024), Data Exploration (Ma et al., 2023), Dialogue act recognition (Qiang et al., 2024); domains like Economics & Finance (Rajpoot et al., 2024; Yu et al., 2023), Politics (Feng et al., 2023), Nutrition & Food (Yang et al., 2024), News (Kuila and Sarkar, 2024); and reasoning types (Huang and Chang, 2023) like ANALOGICAL (Wijesiriwardene et al., 2023) and Multi-hop (Pan et al., 2021) reasoning.
@@
Initially, LMs were relatively weak like GPT-2 (Radford et al., 2019), too large in size like GPT-3 (Brown et al., 2020), expensive like GPT-4 (OpenAI et al., 2024), and/or closed and accessible only via APIs. These pose challenges of accessibility, computational resource limitations (Ding et al., 2024), costs (Jimenez Gutierrez et al., 2022) (GPT-4 Turbo charges \$10/1M input, \$30/1M output tokens), data privacy, inflexibility of customization, and so on. However, there has been recent rise in competitive LMs which are relatively small and openly available. These properties remove the aforementioned limitations, and provide additional benefits like on-device usage, faster inference time, easier compliance and security management.
@@
But do these smaller LMs fit every use case like the SOTA LLMs? They vary a lot in terms of training data, pre-training strategies, and architectural decisions. Further, utilization strategies like zero-shot usage, customizing pre-trained models (e.g., fine-tuning (Mosbach et al., 2023)), using in-context learning (Wei et al., 2022a; Dong et al., 2023), prompt engineering (Brown et al., 2020) can differ with varying need. Writing helpful prompts also requires time and domain expertise. For someone to use these small, open LMs, they need to conduct their analysis with constraints of time, money, computational resources is a complicated task, and identify the LMs fit for their use. Technical reports of some LMs (Team et al., 2024b, c) report performance on different benchmarks, but they are more theoretical than in a practical usage setting. Therefore, there is a need for a comprehensive evaluation of these LMs and determining their capabilities for different use-cases in different practical applications and determine if one can use them instead of SOTA LLMs within tolerable levels of performance degradation.
@@
To bridge this gap, we perform this extensive, in-depth experimental analysis with 10 openly available LMs between 1.7B–11B parameters. We also vary prompt styles by using chat-style task definitions, in-context examples that are closer to how users interact with LMs, and evaluate them across three aspects – task types, application domains and reasoning types, which can be considered as three different ways of categorizing an application area aiming to use LMs. We propose a schema by selecting 12, 12, and 10 entities from each aspect respectively in English language covering a broad range of areas, and group similar entities. For example, in application domains, we group ‘Social Media’ and ‘News’ in ‘Media and Entertainment’. This three-tier structure (aspect, group, entity) allows finding patterns in capabilities of LMs at multiple level, along different aspects.
@@
A recent work (Zhao et al., 2024) dives into efficacy of LoRA (Hu et al., 2022a) fine-tuning of smaller LMs, but uses static prompts, and limits to varying task types only. Further, usage strategy of LMs can significantly vary – fine-tuning (Mosbach et al., 2023), PEFT (Han et al., 2024), multi-stage pipeline based, or directly with/without prompt engineering due to lack of sufficient data, resources, or need for customization. Hence, we consider semantic correctness of outputs as a measure of LMs’ innate ability, and evaluate 5 pre-trained and 5 instruction-tuned (IT) (Ouyang et al., 2022) LMs out-of-the-box with 8 prompt styles. We show that scale can be countered with careful selection of small, open LMs, which can enable matching and outperforming models like GPT-4o-mini, GPT-4o (OpenAI, 2023), DeepSeek-v2 (DeepSeek-AI, 2024) and Gemini-1.5-Pro (Team et al., 2024a) for practical applications. Our proposed framework enables this analysis and identifies patterns in strengths and weaknesses at 3 hierarchical levels.
In this work, we aim to address these research questions: (i) Can small, open LMs compete with large, proprietary LMs in practical usage? (ii) What can be an exhaustive evaluation framework to conduct this analysis? (iii) For different application needs, how do current best small, open LMs perform in comparison, and which LM is the best choice? (iv) What type of prompt style should be used to extract best results from these LMs?
Consistent with Figure 1, we make the following key contributions:
@@
(i) Propose a three-tier evaluation framework to analyze performance of LMs for different task types, application domains and reasoning types.
@@
(ii) Conduct an in-depth experimental analysis of semantic correctness of outputs of 10 open, small LMs in 2B–11B size based on the framework.
(iii) Show that appropriate selection of open, small LMs can lead to outperforming SOTA LLMs like GPT-4o-mini, Gemini-1.5-Pro and DeepSeek-v2, and competing with GPT-4o.
(iv) Compare the performance of LMs with eight prompt styles and recommend the best alternative.
2 Evaluation Framework
@@
We begin with describing our evaluation framework discussing dataset, prompt styles, selection process of aspects, evaluation metrics and experiments.
2.1 Experimental Dataset
@@
We derive our experimental dataset from Super-Natural Instructions (Wang et al., 2022), which is not a single dataset but a meta-dataset constructed by combining many standard NLP datasets. In addition to the source datasets, it also has definition describing a task in chat-style instruction form and many in-context examples (refer Figure 2 for an example) curated by experts. Using datasets from here benefits us by allowing evaluation with various prompt styles and using chat-style instructions – the way users practically interact with LMs. It also provides labels of task type describing nature of a task (eg. question answering, data to text), domain describing the field of the task (eg. history, news), and reasoning type, describing the type of reasoning (if any) needed in the task (eg. multihop, analogical, etc.), which we also use.
@@
2.2 Prompt Styles
We conduct our experiments using multiple prompt styles - including/excluding chat-style task definitions, and with 0, 2, 4, 8 in-context examples for each instance. Examples help LMs (even pre-trained) with in-context learning (Wei et al., 2022a; Dong et al., 2023) without altering their parameters. This is followed by an actual task instance. We select examples from positive examples section of the task. This gives 8 prompt styles per task instance. An example of prompt with definition and 2 examples is given in Figure 2. ‘Input’ and ‘Output’ is used since they are universal for all tasks.
@@ 2.3 Selection of Aspects
From the dataset, we divide each task instance into three aspects – task types, application domains and reasoning types. Since there were many instances for each entity, we filter and rearrange these to create a filtered set for brevity. Our objective was to cover a wide range of application area in each aspect. Therefore, first, we took all the 12 task types in the test set. Among them, for 36 domains and 18 reasoning types, we discarded subsets, very closely similar entities, or ones which didn’t have many examples. For example, there were two domains ‘Computer Science’ and ‘Coding’, so we included only Computer Science as Coding can be considered a subset; among the two types of reasoning called ‘Numerical’ and ‘Quantitative’, we included only Quantitative since they were very similar, and so on. As the number of entities were not too many, we did this manually. We always included the more wider scoped entity when resolving these clashes. After taking a broad enough spectrum in all 3 aspects, we constructed groups in each entity and placed them to create a second-level hierarchy, with similar entities in same groups. Our final structure is shown in Figure 3. Here, Domains is an aspect, Social Sciences and Humanities is a group which contains 4 entities, Economics being one of them. Our intention with this is to provide a structure to this study and cover a broad spectrum of entities. Some of the definitions, specifically for reasoning types, are detailed more in a survey (Guo et al., 2023) and the dataset repository [^2] https://instructions.apps.allenai.org.
This allows analysis at three levels of hierarchy - aspect, group and entity level, which is how we address them in rest of this paper. Some tasks can overlap between entities of same aspect (Kuila and Sarkar, 2024) or different aspects (Keles and Bayraklı, 2024), and some may not belong to any aspect. There are more entities not included here for brevity but listed and evaluated in Appendix B with dataset statistics.
@@ 2.4 Evaluation Metrics
Evaluating LM outputs using n-gram metrics like ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005), etc., have limitations in terms of coherence, consistency, relevance, fluency of those metrics as discussed in recent works (Sai et al., 2021; Xiao et al., 2023) in depth. To address this, we evaluate LM’s knowledge via semantic correctness of outputs using BERTScore (Zhang et al., 2019) recall with roberta-large (Liu et al., 2019) which greatly limits these issues.
Some tasks, like classification, aren’t generation tasks, but we still consider them as one since gives a uniform evaluation paradigm. By aligning outputs using fine-tuning/ICL (Zhao et al., 2023), verbalizers (Hu et al., 2022b), post-processing, labels can be obtained from language outputs.
2.5 Language Models Used
The focus for this work is on open LMs from 1.7–11B parameters for adaptability and computational efficiency. Analysis of pre-trained models, trained for next-word prediction, will give an insight into LMs’ ability and knowledge to perform the tasks. They can either be used directly or adapted/aligned further. IT models will suit out-of-the-box usage on chat-style human-like instructions due to a simple use-case or unavailability of sufficient data/resources to customize the models.
To cover a broad range of SOTA small, open LMs across sizes, families, our experiments utilize Gemma-2B, Gemma-7B (Team et al., 2024b), Llama-3-8B Touvron et al. (2023); AI@Meta (2024), Mistral-7B-v0.3 Jiang et al. (2023), and Falcon-2-11B Almazrouei et al. (2023); TIIUAE (2024). We also take their instruction-tuned (IT) versions (except Falcon-2-11B - not available). But, we omit Mistral-7B pre-trained from main paper’s discussion as its results weren’t competitive, and Gemma-2 series (Team et al., 2024c) since their performance was below Gemma. Model and implementaton details are discussed more in Appendix C,  G. In this paper, suffix "-I" indicates instruction-tuned.
@@ 3 Experiments and Results
We use all the prompt styles with each of the task instance, do a forward pass on the LM, and decode the output using greedy decoding, which is evaluated with available references. We used greedy as it’s reproducible, also other sampling techniques (Holtzman et al., 2020) didn’t give any improvement (refer Appendix E). The following subsections discusses findings.
3.1 Performance Correlation of LMs
@@
One of the hypothesis was that different LMs would perform differently. To demonstrate that, we show the correlation between BERTScore recalls of LM outputs, shown in Figure 4, is low. This shows that their performance with different task types are inherently different, and therefore, selecting the right LM for a usage requirement becomes crucial. To analyze this, we detail their performance in our proposed evaluation framework. For these analyses, we use the best prompt style for that entity of that aspect (refer Appendix D.2 to determine that).
@@ 3.2 Comparison Across Task Types
(d) Task Types; IT Models
(e) Domains; IT Models
(f) Reasonings; IT Models
@@
Most of the pre-trained models perform reasonably well on most tasks. We see that Gemma-2B always and SmolLM-1.7B sometimes perform better than all 7B and 8B models, which is opposite to the general understanding that scale improves performance. So, other design factors are also relevant which contribute to their strengths. Gemma-2B is the best across 50% of the task types, with Falcon-2-11B leading in the remaining, except Word Analogy where SmolLM-1.7B is marginally the best. Considering the scale of the two models, Gemma-2B is a strong choice with resource constraints across all task types, unless Falcon-2-11B is needed purely on performance. Gemma-7B and Llama-3-8B hover below the top two with varying differences. We don’t identify any patterns at group levels here but the difference between the top two models is similar across most tasks.
In IT models, Mistral-7B-I performs best on all task types, with Gemma-2B-I and SmolLM-1.7B-I competing for the second-best. At group level, we find the difference to be smaller for linguistic relationship and generation tasks, but large for semantic & pragmatic analysis tasks. Like their pre-trained variants, Gemma-7B-I and Llama-3-8B-I seldom compete with Gemma-2B-I in some tasks, but never outperforms it. So, Gemma-2B, SmolLM-1.7B-I and Mistral-7B-I can be selected based on performance and resources trade-offs.